OpenStudy (josee):

A point is moving on the graph of xy = 5. When the point is at (4, 5/4 ), its x-coordinate is increasing at a rate of 3 units per second. What is the speed of the y-coordinate at that moment and in which direction is it moving?

OpenStudy (anonymous):

(dx/dt)(y) + x(dy/dt) = 0 (3)(5/4) + (4)(dy/dt) = 0 dy/dt = -15/16

OpenStudy (anonymous):

So it is decreasing at a rate of -15/16 units per second

OpenStudy (josee):

noppe thats not it :/

OpenStudy (josee):

is it decreasing or increasing at y?

OpenStudy (anonymous):

y is decreasing at a rate of 15/16 units per second

OpenStudy (josee):

thx:)